OpenStudy (mallorysipp234):

How do i put this into the calculator and get 6??? It's in the comments.

OpenStudy (mallorysipp234):

\[\log_{9}4 \]

OpenStudy (mallorysipp234):

and \[\log_{36}x=1/2\]

OpenStudy (faiqraees):

log 4 /log 9

OpenStudy (faiqraees):

log(36)x = 1/2 x = 36^1/2

OpenStudy (mallorysipp234):

thank you.